OpenStudy (yanasidlinskiy):

Consider the leading term of the polynomial function. What is the end behavior of the graph? -3x^5 + 9x^4 + 5x^3 + 3 a. The leading term is -3x^5. Since n is odd and a is negative the end behavior is up and up b. The leading term is -3x5. Since n is odd and a is negative the end behavior is down and down c. The leading term is -3x5. Since n is odd and a is negative the end behavior is up and down d. The leading term is -3x5. Since n is odd and a is negative the end behavior is down and up I have C. Is that correct? If not, Please tell me and explain. But explain first!!:D Thanx

OpenStudy (zzr0ck3r):

correct

OpenStudy (johnweldon1993):

Indeed!

OpenStudy (yanasidlinskiy):

Hahha:D Thanx. Cuz I thought I was wrong. My sister was saying it was supposed to be D. I have explained many times!:D Anyways..Thanx again!!:D I need help with another one though.

OpenStudy (zzr0ck3r):

without the - it would be

OpenStudy (yanasidlinskiy):

Are you giz willing to help me with another one?

OpenStudy (zzr0ck3r):

they will always looks like x,-x,x^2,or -x^2

OpenStudy (zzr0ck3r):

what's why we are here

OpenStudy (yanasidlinskiy):

Thanx. Ok. Here is what I need to solve for: \[\sqrt{2x+8} -6 = 4\]

OpenStudy (johnweldon1993):

Step 1...add 6 to both sides

OpenStudy (johnweldon1993):

Step 2... Square both sides in order to eliminate that square root

OpenStudy (yanasidlinskiy):

Wait....You're going to fast Lol..I would add the 6 to the four. Right?

OpenStudy (johnweldon1993):

Correct :)

OpenStudy (yanasidlinskiy):

Ok. So..I have: \[\sqrt{2x+8}=10\]

OpenStudy (johnweldon1993):

So that would leave us with \(\ \sqrt{2x + 8} = 10\) right!

OpenStudy (johnweldon1993):

Now we square both sides \[\large (\sqrt{2x + 8})^2 = (10)^2\]

OpenStudy (johnweldon1993):

Because what that does...is make this happen \[\large 2x + 8 = 100\] It eliminated that square root, because SQUARING a SQUARE ROOT cancels each other out

OpenStudy (yanasidlinskiy):

Ok. Makes sense now;) So now it would be: \[2x=108\] Right? And \[x=54\]

OpenStudy (yanasidlinskiy):

Am I right?

OpenStudy (johnweldon1993):

Ahh not quite....we have \[\large 2x + 8 = 100\] Well since 8 is being ADDED to the 2x...we want to SUBTRACT it to cancel it out...so we need to subtract 8 from both sides of this equation not Add

OpenStudy (yanasidlinskiy):

Oh, wait. Common mistake. Wow. Thanx for noticing it:D \[2x=92\] Divide by 2 and I get 46. That should be correct. Right?

OpenStudy (johnweldon1993):

Indeed :)

OpenStudy (yanasidlinskiy):

Thanx for all your help!!!!:D

OpenStudy (yanasidlinskiy):

It's Algebra 2. Just started this class like Last week.

OpenStudy (johnweldon1993):

Of course :) and yeah it CAN be tricky but once you get it, it seems like elementary. Trust me :P

OpenStudy (yanasidlinskiy):

Hahaha:D Yep! That's really true. Whenever I was In 8th grade I took pre algebra. And I got it down like in a month. But when I started practicing it, Lol, It was like Kindergarten. Then I had to tutor some students!!:D Lol.

OpenStudy (johnweldon1993):

Exactly :) so stick with it, remember to "eliminate" and isolate the variable to solve for it...and you'll be all set :) if not, well I'm always here :D haha

OpenStudy (yanasidlinskiy):

Hahaha:D Yea. I even had to help my older brother with this for college when he was taking his pre-reqs. Lol. Yep! You always explain it sooooooooooooooooooooooooo good:D

OpenStudy (johnweldon1993):

Haha glad to help hun :) alright, I must retire for tonight but if you need anything else feel free to message me again :)